OpenStudy (anonymous):

How do you factor 2m^4-12m^2+16?

OpenStudy (akshay_budhkar):

ok do you know how to solve quadratic equations?

OpenStudy (anonymous):

substitute m^2 by n and then solve it as a quadratic equation.

OpenStudy (akshay_budhkar):

also another hint take 2 common

OpenStudy (akshay_budhkar):

can you do that?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I already know you can pull out a 2 because its the gcf but im not sure what to do after that

OpenStudy (akshay_budhkar):

substitute m^2=n and solve the quadratic equation

OpenStudy (anonymous):

So I would solve 2(m^4-6n+8)?

OpenStudy (akshay_budhkar):

yea but if m^2=n we have m^4=n^2 so you have to solve 2(n^2-6n+8) you can do that right?

OpenStudy (akshay_budhkar):

then resubstitute m after u get the value of n

OpenStudy (anonymous):

ok thanks, i understand the process, ill solve it now. no more help is needed.